OpenStudy (anonymous):

Solve the following systems of equations algebraically: 2y=8-x y-2x=-1 The solution is x=?, y=?

OpenStudy (anonymous):

2y = 8-x we can get x = 8- 2y y-2(8-2y)=-1 y-16+4y= -1 5y = 15 y=3 2*3=8-x 6= 8-x x=2